You are selling tickets for a high school play. Student tickets cost $4 and general admission tickets cost $6. You sell 525 tickets and collect $2876. How many of each type did you sell?
Found 4 solutions by Paul, chripchrip, bodmaris, ikleyn:
Answer by Paul(988)

Let the student tickets be x
Let the geral Admission tickets be y
x+y=525
y=525-x

4x+6y=2876 (subsitute for y)
4x+6(525-x)=2876
4x+3150-6x=2876
-2x=-274
x=137
y=525-137
y=388

Hence, about 137 childern tickets were sold and 388 gernal admission tickets were sold.
Paul.

Answer by chripchrip(6)

let student tickets be x and the general tickets be
525-X
4X+6(525-x)=2876
4x+3150-6x=2876
2x=274
x=137
the nos of student ticket are 137 and the nos of general admission tickets are 388

Answer by bodmaris(2)

I found that it helps to write everything out like this...
Student=$4
General=$6
x=number of student tickets sold
y=number of general admission tickets sold
x+y=525 tickets
4x+6y=2876
THEN you solve it.



Answer by ikleyn(52794)

.
It is a standard ticket problem.

Read the lessons
    - Using systems of equations to solve problems on tickets
    - Three methods for solving standard (typical) problems on tickets
in this site.

From these lessons,  learn on how to solve such problems once for all.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Systems of two linear equations in two unknowns".